A cuboid has a rectangular front face with an are a of 47cm square an a ratio of the length of the cuboid to its width 4:3, whats the volume of the cuboid?
Answer by Alan3354(69443)

A cuboid has a rectangular front face with an are a of 47cm square an a ratio of the length of the cuboid to its width 4:3, whats the volume of the cuboid?
----------------
L*W = 47
L/W = 4/3 --> 3L = 4W --> W = 3L/4
---------
L*(3L/4) = 47
L^2 = 188/3
L =~ 7.916 cm
W =~ 5.937 cm
----------------
Vol = L*W*H = 47*H
No 3rd dimension is given.